# Menta: A Small Language Model for On-Device Mental Health Prediction

### v2

@@ 6 Discussion
While a previous study (jia2025beyond) exploring SLMs in mental health focuses on SLMs in the 2–7B range, our experiments specifically target models strictly smaller than 7B and show that models below about 3.8B parameter scale (e.g., around 1B) fail to meaningfully perform, typically outputting blank or random results in zero- or few-shot settings. This suggests that scaling language models below a certain parameter threshold results in a catastrophic drop in task performance rather than a gradual degradation, indicating the presence of a sharper lower-bound capacity limit. Our work identifies this threshold explicitly within the mental health prediction domain. This observation is consistent with prior findings (sengupta2025compression; subramanian2025small), that performance degradation is not always smooth when size is reduced.
Our findings also align with prior work of MentalQLM (0.5B) (shi2025mentalqlm), which demonstrated strong performance in predicting a range of mental health outcomes including depression, stress, and suicidal ideation. These results highlight the viability of fine-tuned SLMs for mental health prediction, offering a favorable trade-off between computational efficiency and predictive accuracy. Moreover, consistent with recent studies (ren2024learn; yang2024unveiling), we observe that core reasoning and question-answering abilities are preserved after fine-tuning, enabling the resulting models to serve as multi-functional tools with superior performance on mental health tasks.
@@
Our findings were observed in previous work (jia2025beyond) that SLMs often came within 2% of LLMs’ F1 scores in binary classification settings under zero-shot evaluation, highlighting the potential of SLMs as resource-efficient and privacy-conscious alternatives in clinical or low-resource environments. Other prior works have concentrated primarily on LLMs or hybrid models in mental health domains. For instance, Kim et al. (kim2025interpretable) compares zero-shot LLMs and supervised classifiers built on LLM embeddings, showing that while LLMs generalize well in binary depression classification, they struggle in fine-grained severity classification tasks, where our findings collaborate the difficulty in finer labels. Recent surveys (ge2025survey; garg2025rise) highlight the growing promise of SLMs in digital health applications, noting their efficiency, reduced computational demands, and capacity for local deployment, which offer significant advantages over LLMs by being more privacy-preserving and easier to integrate into real-world clinical or mobile settings. Our work complements these surveys by providing empirical evidence for the operational boundary of this trade-off, showing precisely when SLMs transition from viable to non-viable performance levels for such sensitive mental health applications. Also, our compact Menta model demonstrates that the on-device SLM deployment is essential for protecting user privacy, a primary ethical concern in digital psychiatry (patel2018lancet; jin2025applications). In resource-limited clinical settings, particularly in low- and middle-income countries where mental-health workforce shortages persist (moitra2022global), SLMs enable accessible, cost-efficient, and scalable mental-health screening tools. Moreover, smaller models are easier to interpret and audit, aligning with recent calls for transparent and accountable AI in healthcare (doshi2017towards; tonekaboni2019clinicians).
@@
The competitiveness of SLMs in this domain primarily emerges because mental-health language tasks often rely on localized lexical, affective, and syntactic cues, such as expressions of self-reference, emotional polarity, and temporal framing, which can be effectively captured by moderate-sized models without requiring the broad world knowledge encoded in LLMs. As shown in prior work (ji2021mentalbert; shen2018cross; benton2017multi), fine-tuning on domain-specific corpora substantially narrows the performance gap between smaller and larger architectures. Our findings extend this evidence by demonstrating that parameter-efficient fine-tuning (e.g., LoRA) combined with adaptive, weighted loss functions enables SLMs to learn discriminative representations for mental-health prediction at a fraction of the computational cost.
@@
While LLMs generally outperform across a broader range of tasks due to their larger capacity and instruction-following capabilities, SLMs demonstrate competitive and even superior performance on specific tasks, particularly depression and stress prediction tasks. For the remaining tasks, although SLMs show notable improvements, their performance still underperforms compared with LLMs (xu2024mental). This suggests that SLMs can be highly effective when the task structure and linguistic signals are well-aligned with their capacity. However, a critical limitation remains: SLMs struggle with long input sequences due to both token-length constraints and reduced capability in handling extended contextual dependencies (kumar2025large). This restricts their applicability in scenarios requiring nuanced understanding of lengthy or complex user narratives, where LLMs maintain a distinct advantage (lu2024small).
@@
Overall, our findings support a paradigm shift from size-centric scaling to purpose-centric optimization in mental-health AI. As efficiency, transparency, safety and user privacy (zhang2021breaking) become increasingly prioritized in clinical machine learning, the balance between performance and practicality will define the next generation of digital mental-health tools. SLMs address emerging demands in digital mental health for in-situ, low-latency, and privacy-preserving early detection (harari2016using; wang2018tracking), while also aligning with growing calls for transparent and accountable AI in healthcare settings (wies2021digital; smith2023digital). By showing that fine-tuned SLMs with adaptive, cross-dataset strategies can match or surpass LLMs in mental-health outcome prediction, this study provides empirical evidence that smaller, well-aligned models may offer a more sustainable and ethically sound path forward for real-world mental-health monitoring.
@@
What tasks are most suitable for SLMs? In our work, we demonstrate that cross-dataset training is not only feasible but beneficial for predicting depression, stress, and suicidal ideation, when combined with an adaptive loss design. This finding resonates with observations from Yao et al. (benton2017multi), who reported that shared latent representations across related affective tasks can promote positive transfer. However, unlike their multi-task architecture, our approach achieves comparable cross-task gains within a parameter-efficient fine-tuning paradigm (LoRA), thereby reducing computational cost and memory footprint. This distinction highlights that even lightweight fine-tuning frameworks can leverage dataset diversity effectively, provided the optimization objective is properly calibrated.
@@
Also, we found that the fine-tuned model Menta shows better performance in depression and stress detection tasks compared with suicidal risk detection. We found that fine-tuned and non-fine-tuned SLMs, and even larger language models, tend to achieve stronger performance on depression and stress detection tasks compared to suicide risk classification. For example, a systematic review (bauer2024using) further notes that language models often misclassify nuanced expressions of suicidality due to class imbalance and the reliance on contextual or temporal dependencies, which are not captured in single posts. Our findings also collaborate with another study (lamichhane2023evaluation), which reported F1‑scores of 0.86 for depression detection and only 0.37 for suicidality detection on social media data. These findings reflect that depression and stress often manifest with clearer linguistic patterns and more abundant data, whereas suicidality may be rarer, more context dependent, and linguistically subtler, making it harder for models to learn reliably. As a consequence, SLMs (and even larger models) may be better suited for screening tasks (e.g., detecting depression or stress) than for nuanced stratification of suicide risk without additional multimodal or temporal data.
@@
While our evaluation used diverse datasets covering depression, stress, and suicidal ideation, all inputs were textual social media posts. This focus excludes multimodal features such as linguistic style dynamics, user metadata, or interaction patterns, that often play a critical role in understanding real-world mental health contexts. Incorporating multimodal signals could improve robustness and ecological validity in future work (yazdavar2020multimodal; khoo2024machine). Also, one key limitation of this study lies in the restricted diversity of datasets and mental health tasks explored. Our evaluation primarily focused on three task categories depression, stress, and suicidal ideation, each derived from publicly available social media datasets. While these benchmarks are widely adopted in affective computing, they still represent a narrow subset of the psychological spectrum. Real-world mental health expressions often encompass more nuanced and overlapping conditions such as anxiety, bipolar disorder, and emotional dysregulation, which are underrepresented in our current corpus (cao2024machine). Additionally, all analyses were conducted on publicly available, anonymized datasets rather than in clinical environments. While this design ensures privacy and reproducibility, it limits conclusions about clinical safety, ethical compliance, and longitudinal reliability (seyedsalehi2024suicide). Future directions include the test of federated or on-device fine-tuned SLMs in clinical settings. Deploying such models as assistive rather than diagnostic tools, under clinician supervision, could provide a safe bridge between algorithmic insights and responsible mental health support.

### v3

@@ 6 Discussion
While a previous study (Jia et al., 2025) exploring SLMs in mental health focuses on SLMs in the 2–7B range, our experiments specifically target models strictly smaller than 7B and show that models below about 3.8B parameter scale (e.g., around 1B) fail to meaningfully perform, typically outputting blank or random results in zero- or few-shot settings. This suggests that scaling language models below a certain parameter threshold results in a catastrophic drop in task performance rather than a gradual degradation, indicating the presence of a sharper lower-bound capacity limit. Our work identifies this threshold explicitly within the mental health prediction domain. This observation is consistent with prior findings (Sengupta et al., 2025; Subramanian et al., 2025), that performance degradation is not always smooth when size is reduced.
Our findings also align with prior work of MentalQLM (0.5B) (Shi et al., 2025), which demonstrated strong performance in predicting a range of mental health outcomes including depression, stress, and suicidal ideation. These results highlight the viability of fine-tuned SLMs for mental health prediction, offering a favorable trade-off between computational efficiency and predictive accuracy. Moreover, consistent with recent studies (Ren et al., 2024; Yang et al., 2024a), we observe that core reasoning and question-answering abilities are preserved after fine-tuning, enabling the resulting models to serve as multi-functional tools with superior performance on mental health tasks.
@@
Our findings were observed in previous work (Jia et al., 2025) that SLMs often came within 2% of LLMs’ F1 scores in binary classification settings under zero-shot evaluation, highlighting the potential of SLMs as resource-efficient and privacy-conscious alternatives in clinical or low-resource environments. Other prior works have concentrated primarily on LLMs or hybrid models in mental health domains. For instance, Kim et al. (Kim et al., 2025) compares zero-shot LLMs and supervised classifiers built on LLM embeddings, showing that while LLMs generalize well in binary depression classification, they struggle in fine-grained severity classification tasks, where our findings collaborate the difficulty in finer labels. Recent surveys (Ge et al., 2025; Garg et al., 2025) highlight the growing promise of SLMs in digital health applications, noting their efficiency, reduced computational demands, and capacity for local deployment, which offer significant advantages over LLMs by being more privacy-preserving and easier to integrate into real-world clinical or mobile settings. Our work complements these surveys by providing empirical evidence for the operational boundary of this trade-off, showing precisely when SLMs transition from viable to non-viable performance levels for such sensitive mental health applications. Also, our compact Menta model demonstrates that the on-device SLM deployment is essential for protecting user privacy, a primary ethical concern in digital psychiatry (Patel et al., 2018; Jin et al., 2025). In resource-limited clinical settings, particularly in low- and middle-income countries where mental-health workforce shortages persist (Moitra et al., 2022), SLMs enable accessible, cost-efficient, and scalable mental-health screening tools. Moreover, smaller models are easier to interpret and audit, aligning with recent calls for transparent and accountable AI in healthcare (Doshi-Velez & Kim, 2017; Tonekaboni et al., 2019).
@@
The competitiveness of SLMs in this domain primarily emerges because mental-health language tasks often rely on localized lexical, affective, and syntactic cues, such as expressions of self-reference, emotional polarity, and temporal framing, which can be effectively captured by moderate-sized models without requiring the broad world knowledge encoded in LLMs. As shown in prior work (Ji et al., 2021; Shen et al., 2018; Benton et al., 2017), fine-tuning on domain-specific corpora substantially narrows the performance gap between smaller and larger architectures. Our findings extend this evidence by demonstrating that parameter-efficient fine-tuning (e.g., LoRA) combined with adaptive, weighted loss functions enables SLMs to learn discriminative representations for mental-health prediction at a fraction of the computational cost.
@@
While LLMs generally outperform across a broader range of tasks due to their larger capacity and instruction-following capabilities, SLMs demonstrate competitive and even superior performance on specific tasks, particularly depression and stress prediction tasks. For the remaining tasks, although SLMs show notable improvements, their performance still underperforms compared with LLMs (Xu et al., 2024). This suggests that SLMs can be highly effective when the task structure and linguistic signals are well-aligned with their capacity. However, a critical limitation remains: SLMs struggle with long input sequences due to both token-length constraints and reduced capability in handling extended contextual dependencies (Kumar, 2025). This restricts their applicability in scenarios requiring nuanced understanding of lengthy or complex user narratives, where LLMs maintain a distinct advantage (Lu et al., 2024).
@@
Overall, our findings support a paradigm shift from size-centric scaling to purpose-centric optimization in mental-health AI. As efficiency, transparency, safety and user privacy (Zhang et al., 2021) become increasingly prioritized in clinical machine learning, the balance between performance and practicality will define the next generation of digital mental-health tools. SLMs address emerging demands in digital mental health for in-situ, low-latency, and privacy-preserving early detection (Harari et al., 2016; Wang et al., 2018), while also aligning with growing calls for transparent and accountable AI in healthcare settings (Wies et al., 2021; Smith et al., 2023). By showing that fine-tuned SLMs with adaptive, cross-dataset strategies can match or surpass LLMs in mental-health outcome prediction, this study provides empirical evidence that smaller, well-aligned models may offer a more sustainable and ethically sound path forward for real-world mental-health monitoring.
@@
What tasks are most suitable for SLMs? In our work, we demonstrate that cross-dataset training is not only feasible but beneficial for predicting depression, stress, and suicidal ideation, when combined with an adaptive loss design. This finding resonates with observations from Yao et al. (Benton et al., 2017), who reported that shared latent representations across related affective tasks can promote positive transfer. However, unlike their multi-task architecture, our approach achieves comparable cross-task gains within a parameter-efficient fine-tuning paradigm (LoRA), thereby reducing computational cost and memory footprint. This distinction highlights that even lightweight fine-tuning frameworks can leverage dataset diversity effectively, provided the optimization objective is properly calibrated.
@@
Also, we found that the fine-tuned model Menta shows better performance in depression and stress detection tasks compared with suicidal risk detection. We found that fine-tuned and non-fine-tuned SLMs, and even larger language models, tend to achieve stronger performance on depression and stress detection tasks compared to suicide risk classification. For example, a systematic review (Bauer et al., 2024) further notes that language models often misclassify nuanced expressions of suicidality due to class imbalance and the reliance on contextual or temporal dependencies, which are not captured in single posts. Our findings also collaborate with another study (Lamichhane, 2023), which reported F1‑scores of 0.86 for depression detection and only 0.37 for suicidality detection on social media data. These findings reflect that depression and stress often manifest with clearer linguistic patterns and more abundant data, whereas suicidality may be rarer, more context dependent, and linguistically subtler, making it harder for models to learn reliably. As a consequence, SLMs (and even larger models) may be better suited for screening tasks (e.g., detecting depression or stress) than for nuanced stratification of suicide risk without additional multimodal or temporal data.
@@
While our evaluation used diverse datasets covering depression, stress, and suicidal ideation, all inputs were textual social media posts. This focus excludes multimodal features such as linguistic style dynamics, user metadata, or interaction patterns, that often play a critical role in understanding real-world mental health contexts. Incorporating multimodal signals could improve robustness and ecological validity in future work (Yazdavar et al., 2020; Khoo et al., 2024). Also, one key limitation of this study lies in the restricted diversity of datasets and mental health tasks explored. Our evaluation primarily focused on three task categories depression, stress, and suicidal ideation, each derived from publicly available social media datasets. While these benchmarks are widely adopted in affective computing, they still represent a narrow subset of the psychological spectrum. Real-world mental health expressions often encompass more nuanced and overlapping conditions such as anxiety, bipolar disorder, and emotional dysregulation, which are underrepresented in our current corpus (Cao et al., 2024). Additionally, all analyses were conducted on publicly available, anonymized datasets rather than in clinical environments. While this design ensures privacy and reproducibility, it limits conclusions about clinical safety, ethical compliance, and longitudinal reliability (Seyedsalehi & Fazel, 2024). Future directions include the test of federated or on-device fine-tuned SLMs in clinical settings. Deploying such models as assistive rather than diagnostic tools, under clinician supervision, could provide a safe bridge between algorithmic insights and responsible mental health support.